# FORECASTOR – II. Simulating Galaxy Surveys with the Cosmological Advanced Survey Telescope for Optical and UV Research

### v1

@@ Abstract
The Cosmological Advanced Survey Telescope for Optical and UV Research (CASTOR) is a planned flagship space telescope, covering the blue-optical and UV part of the spectrum. Here we introduce the CASTOR image simulator, a Python GalSim package-based script capable of generating mock CASTOR images from an input catalogue. We generate example images from the CASTOR Wide, Deep, and Ultra-Deep surveys using simulated light-cones from the Santa Cruz Semi-Analytic Model. We make predictions for the performance of these surveys by comparing galaxies that are extracted from each image using Source Extractor to the input catalogue. We find that the Wide, Deep, and Ultra-Deep surveys will be complete to , 29 and 30 mag, respectively, in the UV, u, and g filters, with the UV-split and u-split filters reaching a shallower depth. With a large area of deg, the Wide survey will detect hundreds of millions of galaxies out to , mostly with . The Ultra-Deep survey will probe to , detecting a large fraction of galaxies. These powerful samples will enable precision measurements of the distribution of star formation in the cosmic web, connecting the growth of stellar mass to the assembly of dark matter halos over two thirds of the history of the Universe, and other core goals of CASTOR’s legacy surveys. These image simulations and the tools developed to generate them will be a vital planning tool to estimate CASTOR’s performance and iterate the telescope and survey designs prior to launch.
1 Introduction
@@
The Cosmological Advanced Survey Telescope for Optical and UV Research (CASTOR) is a planned 1m-class UV/optical telescope led jointly by the National Research Council of Canada and the Canadian Space Agency. CASTOR will offer simultaneous UV, u, and g-band imaging with angular resolution over a 0.25 deg field of view, with a filter wheel including a bandpass filter and a grism, as well as UV multi-object spectroscopy in a parallel field (Côté et al., 2019a, b). Enabled by these unique capabilities, CASTOR’s planned surveys will explore the evolution of galaxies at Cosmic Noon, probe the origins of cosmic explosions, unlock a new era of precision cosmology, test the habitability of distant worlds, and unveil the nature of everything from the composition of icy bodies in the outer solar system to the masses of distant quasars.
@@ 2 Planned CASTOR Extragalactic Surveys
Within the context of galaxy surveys, three key legacy surveys planned for CASTOR are the Wide, Deep and Ultra-Deep surveys, which together comprised a tiered “wedding-cake” approach to CASTOR’s galaxy evolution science (a fourth survey, the CASTOR Nearby Galaxies Survey, is beyond the scope of this work). The exposure times and survey areas of these surveys are summarized in Table 1. The Wide Survey will image the 2227 deg Roman High Latitude Wide Area Survey (HLS) field to a point source depth of 27–28 mag; the Deep Survey will image 83 deg in six contiguous regions overlapping the Rubin/Euclid Deep fields, reaching a point source depth of mag; and finally, the Ultra-Deep survey will span four CASTOR pointings, totalling 1 deg, to a depth of –31 mag (in this paper we predict completeness limits as given in Table 2). Together, these surveys will trace the distribution of star formation within the cosmic web, enable precision photometric and spectroscopic redshift measurements, connect the growth of stellar mass over cosmic time to the assembly of dark matter halos, probe the outer reaches of massive, quiescent galaxies, identify a large sample of faint, post-starburst galaxies, and extend star formation rate and morphology measurements of dwarf galaxies out to high redshifts.
@@
Survey | Exp. Time (s) | Area (deg) | |
@@ 3.1 GalSim
GalSim (Rowe et al., 2015) is an open-source Python package that allows the user to simulate images of astrophysical objects (stars and galaxies) assuming basic radial emission profiles (e.g. point source and Sérsic). GalSim is a widely used and adaptable package for producing mock galaxy images, which has been used to make detailed predictions for a range of upcoming instruments such as Euclid (e.g. Euclid Collaboration et al., 2022a, 2023), Roman (e.g. Troxel et al., 2021, 2023; Drakos et al., 2022) and the Rubin Observatory (e.g. Sánchez et al., 2020; LSST Dark Energy Science Collaboration et al., 2021). Although the GalSim images assume basic profiles, unlike for example Marshall et al. (2022) and Fortuni et al. (2023) that use hydrodynamical simulations to model realistic galaxy structure, this is a simple, robust, and reliable strategy for estimating the overall performance of a telescope. In this work, the CASTOR image simulator uses GalSim to generate mock CASTOR images based on current design specifications of the telescope.
@@
Necessary input files include the point spread function (PSF) model and the bandpass files for the UV, UV-split, u, u-split, and g filters, provided by the CASTOR collaboration (Cheng et al., 2024). For the PSF model we use realistic models provided by CASTOR’s primary Phase 0 contractor Honeywell Aerospace (private communication), which simulate multiple realisations of the PSF at different detector locations. We take the median PSF from the various realisations, and use the PSF from the centre of the detector, assuming that any spatial variations are negligible. We also convolve the model PSF with a Gaussian of to account for spacecraft jitter. The image simulator uses the key CASTOR imaging specifications valid as of the time of writing: 1.0 m aperture diameter, pixel scale of /pixel, 1.0 e-/ADU gain, and a by (0.25 deg) field of view (FOV). These parameters can easily be modified in the script to simulate other telescopes.
Four noise terms are applied to the generated images: sky noise, read noise, dark current, and shot noise. The sky background, including a zodiacal, earthshine, and geocoronal emission line contribution, is computed from Hubble Space Telescope (HST) continuum data files (STScI Development Team, 2013). We implement the “high” earthshine model used in the HST ETC, defined as the earthshine at 38° from the limb (Medallon et al., 2023). For the zodiacal light we use the template used in the HST ETC from Diaz (2015), normalized to a Johnson V magnitude of 22.7 mag corresponding to the “medium” zodiacal level; we investigate how the assumed zodiacal light level affects our completeness limits in Appendix A. Our sky-noise function outputs a single sky background flux value based on the integrated product of the chosen filter response function and this background (earthshine plus zodiacal) continuum spectrum. Airglow spikes as seen through the filter response are then added. We assume the “average” airglow estimate, with the [OIII] line at 2471Å having an intensity of erg/s/cm/arcsec (Medallon et al., 2023).
The sky background flux is converted from its output value in units of photons/s/cm/pixel to its value in units of ADU/pixel using
@@
To create astrophysical images based on these telescope parameters, we require an object catalogue. In this work, we use simulated light-cones, which are described in Section 3.2.
@@ 3.2 Simulated Light-Cones
The image simulator can take a light-cone catalogue as input and generate a corresponding field as seen by CASTOR. The input catalogue must contain the positions, Sérsic properties, and fluxes of each object as required by GalSim. In this work, we use a set of five realizations of 2-deg lightcones constructed using the Santa Cruz semi-analytic model (SAM) for galaxy formation based on the same pipeline presented in Yung et al. (2023). The custom version of lightcones made for this work include additional photometry computed for CASTOR filters. This simulated lightcone is and contains galaxies over the redshift range and rest-frame UV magnitude range mag.
The dark matter halos in these lightcones are extracted from the SMDPL simulation from the MultiDark suite (Klypin et al., 2016), which spans (400 Mpc h) and resolves halos down to M. These are projected onto a mock observed field using the lightcone package that is released as part of the UniverseMachine package (Behroozi et al., 2019). A total of five realizations of a 2-deg field are generated by sampling different parts of the cosmological simulations, providing a total of deg of simulated sky coverage. The dark matter halos in the lightcones preserve their relative distances and spatial distributions from the underlying cosmological simulation and provide redshift evolution along the line of sight that mimics the way we observe the real Universe. For halos in the lightcone, Monte Carlo merger trees are generated with the extended Press-Schechter formalism (e.g. Somerville & Kolatt, 1999) and the Santa Cruz SAM is used to track the evolution of galaxies therein.
The Santa Cruz SAM consists of a set of carefully curated recipes for galaxy formation and evolution that are either derived analytically or from observations and hydrodynamic simulations (Somerville & Primack, 1999; Somerville et al., 2015, 2021). These standard physical processes include gas cooling and accretion, multi-phase ISM gas partitioning, H-based star formation, stellar feedback, galaxy mergers, chemical evolution, black hole growth, and active galactic nuclei (AGN) feedback. The models have been shown to well-reproduce the observed one-point distribution functions of , , and star formation rate (Yung et al., 2019a, b), observational constraints on the intergalactic medium (IGM) reionization history (Yung et al., 2020b, a), AGN populations (Yung et al., 2021), as well as two-point auto-correlation functions from (Yung et al., 2022). We refer the reader to the schematic flow chart in Yung et al. (2022, fig. 1) for a comprehensive illustration of the lightcone construction pipeline and the internal workflow of the Santa Cruz SAM, and Table A1 in Yung et al. (2023) for the full set of quantities available in the mock lightcone. Here, we provide a concise summary of the model components that are most relevant to this work.
@@
This catalogue includes five realizations of 2 deg fields, each of which can contain six non-overlapping CASTOR pointings, making a total of 30 independent pointings. In this work we consider only one of the 2-deg realizations, with 6 independent CASTOR pointings, which is a sufficiently large area for our study. The 2-deg SAM light-cones provide a wide range of properties for each simulated object in the field, including but not limited to observable properties such as magnitudes in specific bands, redshift, celestial coordinates, and disk and bulge Sérsic radii, and physical properties from the simulation such as average star formation rates, and stellar masses. From the light-cone catalogue, for each galaxy we use the properties: celestial coordinates (RA and Dec), intrinsic magnitude of the disk and bulge separately, combined dust-attenuated magnitude of both the disk and bulge, the disk Sérsic radius, bulge radius, minor-to-major axis ratio as specified by where is the inclination angle, and redshift. We randomly assign each galaxy with a position angle. We assume that the disk component of each galaxy has a Sérsic index , while bulges have .
@@
The 2-deg SAM catalogue provides the total galaxy dust-attenuated magnitudes, however it does not provide dust-attenuated magnitudes for the galaxy disk and bulge separately. Thus we estimate the separate dust-attenuated magnitudes by assuming that the bulge flux to total flux, and disk flux to total flux ratios, are equal for intrinsic and dust-attenuated fluxes:
where all variables are fluxes, the D subscript signifies ‘dust’ flux values, and the ND subscript signifies ‘no dust’ flux values.
@@
The flux density is converted to a flux in photons/s/cm by integrating the product of the chosen filter response and the flux density over the spectral range. The flux is then converted to ADU/pixel using Equation 1.
@@
Each galaxy is convolved with the CASTOR PSF for the chosen filter using the GalSim function Convolve. The single galaxy image is then added to the full field image at the position corresponding to the galaxy’s coordinates. After cycling through all galaxies in the catalogue, noise is added to the full field image as described in Section 3.1. The image is saved as a FITS file. Using a 2 deg field from the SAM catalogue, each full field image contains over 1.5 million galaxies.
@@ 4.1 CASTOR survey images
Here we present example CASTOR images created using our image simulator, for each of the planned CASTOR surveys: Wide, Deep, and Ultra-Deep. For each exposure time and filter combination, we create six images of various CASTOR FOVs within the 2-deg SAM catalogue. This allows us to test our codes and predictions with more galaxies, as well as giving estimates for the effect of cosmic variance on the number of galaxies expected in each image. In this section we show only one of these FOVs as an example, although we use all of these in our quantitative analysis in Section 5. We note that our images are made only using a galaxy catalogue, and include no foreground stars. Thus we cannot investigate issues such as star–galaxy blending, where some of the detectable galaxies in our images would, in reality, be obscured and undetectable because of overlapping foreground stars. However, this allows us to more easily extract sources and perform a comparison to our input galaxy catalogue, which is adequate for our galaxy-focused study. With the high resolution of CASTOR, we expect that this will have a negligible impact on the completeness limits that we estimate in this work.
@@ 4.2 Comparison to GALEX and Rubin
To simulate GALEX images for a comparison to CASTOR, we use its telescope parameters diameter m, pixel scale /pixel, a gain of 1.0 e-/ADU, with no dark current or read noise. We simulate images in the NUV filter, assuming a PSF FWHM of , which we assume is a Gaussian in shape (Morrissey et al., 2005), and a sky background level of 26.5 mag/arcsec (Martin et al., 2005). For exposure times, we use those from the All-sky Imaging Survey (AIS), 0.1ks, Medium Imaging Survey (MIS), 1.5ks, Deep Imaging Survey (DIS), 30ks, and the Ultra-Deep Imaging Survey (UDIS), 200ks (Martin et al., 2005). The resulting images are shown in Figure 3. These clearly show the improvement in UV-imaging that CASTOR will have over GALEX, with significantly improved depth and resolution. The GALEX resolution was in the NUV, compared to CASTOR’s planned . This 30-fold improvement in resolution will mitigate the problem of source blending in GALEX imaging, and significantly improve our ability to identify UV sources and accurately measure their properties — both in the local universe and across cosmic timescales (Martin et al., 2005; Noeske et al., 2007; Salim et al., 2007; Gil de Paz et al., 2007; Thilker et al., 2007).
@@
To make simulated Rubin images, we use a diameter m, a pixel scale of /pixel, dark current of 0.02 e-/pixel/s, a read noise of 9 e-/pixel, and a gain of 1.0 e-/ADU (O’Connor, 2019). We simulate images in the g-band, with a sky background of 22.26 mag/arcsec (Ivezić et al., 2019). We assume that the PSF is the convolution of two Gausssians, one with FWHM and one with , each with the same flux, to be consistent with Imsim, the LSST simulation software package111https://lsstdesc.org/imSim. We simulate exposure times of 30s, corresponding to a single visit, 300s, the 1 year depth, 3000s, the 10 year depth, and 39000s, the final depth in each of LSST’s four Deep Drilling Fields (see Côté et al. 2024, in preparation). The resulting images are shown in Figure 3. While the improvement of the CASTOR images is not as dramatic as for GALEX, the zoomed-in panels still clearly show the effect of the increased resolution of CASTOR over LSST. With LSST’s FWHM due to atmospheric seeing, the space-based CASTOR offers a factor of improvement in spatial resolution.

### v2

@@ Abstract
The Cosmological Advanced Survey Telescope for Optical and UV Research (CASTOR) is a planned flagship space telescope, covering the blue-optical and UV part of the spectrum. Here we introduce the CASTOR image simulator, a Python GalSim package-based script capable of generating mock CASTOR images from an input catalogue. We generate example images from the CASTOR Wide, Deep, and Ultra-Deep surveys using simulated light-cones from the Santa Cruz Semi-Analytic Model. We make predictions for the performance of these surveys by comparing galaxies that are extracted from each image using Source Extractor to the input catalogue. We find that the Wide, Deep, and Ultra-Deep surveys will be 75% complete for point sources down to , 29 and 30 mag, respectively, in the UV, u, and g filters, with the UV-split and u-split filters reaching a shallower depth. With a large area of deg2, the Wide survey will detect hundreds of millions of galaxies out to , mostly with . The Ultra-Deep survey will probe to , detecting galaxies with . These galaxy samples will enable precision measurements of the distribution of star formation in the cosmic web, connecting the growth of stellar mass to the assembly of dark matter halos over two thirds of the history of the Universe, and other core goals of CASTOR’s legacy surveys. These image simulations and the tools developed to generate them will be a vital planning tool to estimate CASTOR’s performance and iterate the telescope and survey designs prior to launch.
keywords:
Space telescopes(1547); Ultraviolet astronomy(1736); Ultraviolet surveys(1742); Galaxy evolution(594); Star formation(1569); Astronomical simulations(1857)1 Introduction
@@
The Cosmological Advanced Survey Telescope for Optical and UV Research (CASTOR) is a planned 1m-class UV/optical telescope led jointly by the National Research Council of Canada and the Canadian Space Agency, in bilateral partnership with the United Kingdom Space Agency. CASTOR will offer simultaneous UV, u, and g-band imaging with angular resolution over a 0.25 deg2 field of view, with a filter wheel including a bandpass filter and a grism, as well as UV multi-object spectroscopy in a parallel field (Côté et al., 2019a, b). Enabled by these unique capabilities, CASTOR’s planned surveys will explore the evolution of galaxies at Cosmic Noon, probe the origins of cosmic explosions, unlock a new era of precision cosmology, test the habitability of distant worlds, and unveil the nature of everything from the composition of icy bodies in the outer solar system to the masses of distant quasars.
@@ 2 Planned CASTOR Extragalactic Surveys
Within the context of galaxy surveys, three key legacy surveys planned for CASTOR are the Wide, Deep and Ultra-Deep surveys, which together comprised a tiered “wedding-cake” approach to CASTOR’s galaxy evolution science (a fourth survey, the CASTOR Nearby Galaxies Survey, is beyond the scope of this work). The exposure times and survey areas of these surveys are summarized in Table 1. The Wide Survey will image the 2227 deg2 Roman High Latitude Wide Area Survey (HLS) field to a point source depth of 27–28 mag; the Deep Survey will image 83 deg2 in six contiguous regions overlapping the Rubin/Euclid Deep fields, reaching a point source depth of mag; and finally, the Ultra-Deep survey will span four CASTOR pointings, totalling 1 deg2, to a depth of –31 mag (in this paper we predict completeness limits as given in Table 2). Together, these surveys will trace the distribution of star formation within the cosmic web, enable precision photometric and spectroscopic redshift measurements, connect the growth of stellar mass over cosmic time to the assembly of dark matter halos, probe the outer reaches of massive, quiescent galaxies, identify a large sample of faint, post-starburst galaxies, and extend star formation rate and morphology measurements of dwarf galaxies out to high redshifts.
@@
Survey | Exp. Time (s) | Area (deg2) | |
@@ 3.1 GalSim
GalSim (Rowe et al., 2015) is an open-source Python package that allows the user to simulate images of astrophysical objects (stars and galaxies) assuming basic radial emission profiles (e.g. point source and Sérsic) or allowing for arbitrary morphologies with input 2D brightness distributions. GalSim is a widely used and adaptable package for producing mock galaxy images, which has been used to make detailed predictions for a range of upcoming instruments such as Euclid (e.g. Euclid Collaboration et al., 2022a, 2023), Roman (e.g. Troxel et al., 2021, 2023; Drakos et al., 2022) and the Rubin Observatory (e.g. Sánchez et al., 2020; LSST Dark Energy Science Collaboration et al., 2021). In this work, the CASTOR image simulator uses GalSim to generate mock CASTOR images based on current design specifications of the telescope.
@@
Necessary input files include the point spread function (PSF) model and the bandpass files for the UV, UV-split, u, u-split, and g filters, provided by the CASTOR collaboration (Cheng et al., 2024). For the PSF model we use realistic models provided by CASTOR’s primary Phase 0 contractor Honeywell Aerospace (private communication), which simulate multiple realisations of the PSF at different detector locations. We take the median PSF from the various realisations, and use the PSF from the centre of the detector, assuming that any spatial variations are negligible. We also convolve the model PSF with a Gaussian of to account for spacecraft jitter. The image simulator uses the key CASTOR imaging specifications valid as of the time of writing: 1.0 m aperture diameter, pixel scale of /pixel, 1.0 e-/ADU gain, and a by (0.25 deg2) field of view (FOV). These parameters can easily be modified in the script to simulate other telescopes.
Four noise terms are applied to the generated images: sky noise, read noise, dark current, and shot noise. The sky background, including a zodiacal, earthshine, and geocoronal emission line contribution, is computed from Hubble Space Telescope (HST) continuum data files (STScI Development Team, 2013). We implement the “high” earthshine model used in the HST ETC, defined as the earthshine at 38° from the limb (Medallon et al., 2023). For the zodiacal light we use the template used in the HST ETC from Diaz (2015), normalized to a Johnson V magnitude of 22.7 mag corresponding to the “medium” zodiacal level; we investigate how the assumed zodiacal light level affects our completeness limits in Appendix A. Our sky-noise function outputs a single sky background flux value based on the integrated product of the chosen filter response function and this background (earthshine plus zodiacal) continuum spectrum. Airglow spikes as seen through the filter response are then added. We assume the “average” airglow estimate, with the [OIII] line at 2471Å having an intensity of erg/s/cm2/arcsec2 (Medallon et al., 2023).
The sky background flux is converted from its output value in units of photons/s/cm2/pixel to its value in units of ADU/pixel using
@@
To create astrophysical images based on these telescope parameters, we require an object catalogue. In this work, we use simulated light-cones, which are described in Section 3.2. The model that produces the light-cones assumes that each galaxy can be described by two Sérsic profiles, representing a disk and bulge component. This is a much simpler model than, for example, in Marshall et al. (2022) and Fortuni et al. (2023) where hydrodynamical simulations are used to model realistic galaxy structure in their mock imaging. While GalSim could simulate images using more detailed input galaxy profiles, this is significantly more computationally expensive than assuming a basic profile, and would be prohibitive for simulating the very large CASTOR FOVs. Assuming basic Sérsic profiles is a simple, robust, and reliable strategy for estimating the overall performance of a telescope.
@@ 3.2 Simulated Light-Cones
Our image simulator can take a light-cone catalogue as input and generate a corresponding field as seen by CASTOR. The input catalogue must contain the positions, Sérsic properties, and fluxes of each object as required by GalSim. In this work, we use a set of five realizations of 2-deg2 lightcones constructed using the Santa Cruz semi-analytic model (SAM) for galaxy formation based on the same pipeline presented in Yung et al. (2023). The custom version of lightcones made for this work include additional photometry computed for CASTOR filters. This simulated lightcone is and contains galaxies over the redshift range and rest-frame UV magnitude range mag.
The dark matter halos in these lightcones are extracted from the SMDPL simulation from the MultiDark suite (Klypin et al., 2016), which spans (400 Mpc h-1)3 and resolves halos down to M⊙. These are projected onto a mock observed field using the lightcone package that is released as part of the UniverseMachine package (Behroozi et al., 2019). A total of five realizations of a 2-deg2 field are generated by sampling different parts of the cosmological simulations, providing a total of deg2 of simulated sky coverage. The dark matter halos in the lightcones preserve their relative distances and spatial distributions from the underlying cosmological simulation and provide redshift evolution along the line of sight that mimics the way we observe the real Universe. For halos in the lightcone, Monte Carlo merger trees are generated with the extended Press-Schechter formalism (e.g. Somerville & Kolatt, 1999) and the Santa Cruz SAM is used to track the evolution of galaxies therein.
The Santa Cruz SAM consists of a set of carefully curated recipes for galaxy formation and evolution that are either derived analytically or from observations and hydrodynamic simulations (Somerville & Primack, 1999; Somerville et al., 2015, 2021). These standard physical processes include gas cooling and accretion, multi-phase ISM gas partitioning, H2-based star formation, stellar feedback, galaxy mergers, chemical evolution, black hole growth, and active galactic nuclei (AGN) feedback. The models have been shown to well-reproduce the observed one-point distribution functions of , , and star formation rate (Yung et al., 2019a, b), observational constraints on the intergalactic medium (IGM) reionization history (Yung et al., 2020a, b), AGN populations (Yung et al., 2021), as well as two-point auto-correlation functions from (Yung et al., 2022). We refer the reader to the schematic flow chart in Yung et al. (2022, fig. 1) for a comprehensive illustration of the lightcone construction pipeline and the internal workflow of the Santa Cruz SAM, and Table A1 in Yung et al. (2023) for the full set of quantities available in the mock lightcone. Here, we provide a concise summary of the model components that are most relevant to this work.
@@
This catalogue includes five realizations of 2 deg2 fields, each of which can contain six non-overlapping CASTOR pointings, making a total of 30 independent pointings. In this work we consider only one of the 2-deg2 realizations, with 6 independent CASTOR pointings, which is a sufficiently large area for our study. The 2-deg2 SAM light-cones provide a wide range of properties for each simulated object in the field, including but not limited to observable properties such as magnitudes in specific bands, redshift, celestial coordinates, and disk and bulge Sérsic radii, and physical properties from the simulation such as average star formation rates, and stellar masses. From the light-cone catalogue, for each galaxy we use the properties: celestial coordinates (RA and Dec), intrinsic magnitude of the disk and bulge separately, combined dust-attenuated magnitude of both the disk and bulge, the disk Sérsic radius, bulge radius, minor-to-major axis ratio as specified by where is the inclination angle, and redshift. We randomly assign each galaxy with a position angle. We assume that the disk component of each galaxy has a Sérsic index , while bulges have .
@@
The 2-deg2 SAM catalogue provides the total galaxy dust-attenuated magnitudes, however it does not provide dust-attenuated magnitudes for the galaxy disk and bulge separately. Thus we estimate the separate dust-attenuated magnitudes by assuming that the bulge flux to total flux, and disk flux to total flux ratios, are equal for intrinsic and dust-attenuated fluxes:
where all variables are fluxes, the D subscript signifies ‘dust’ flux values, and the ND subscript signifies ‘no dust’ flux values. In reality, the bulge and disk components would have different levels of dust attenuation, however our simple assumption is the most reasonable estimate given the information that is available. As we only use the SAM to understand basic source detectability in this work, we do not expect that this approximation will have any significant impact on our results.
@@
The flux density is converted to a flux in photons/s/cm2 by integrating the product of the chosen filter response and the flux density over the spectral range. The flux is then converted to ADU/pixel using Equation 1.
@@
Each galaxy is convolved with the CASTOR PSF for the chosen filter using the GalSim function Convolve. The single galaxy image is then added to the full field image at the position corresponding to the galaxy’s coordinates. After cycling through all galaxies in the catalogue, noise is added to the full field image as described in Section 3.1. The image is saved as a FITS file. Using a 2 deg2 field from the SAM catalogue, each full field image contains over 1.5 million galaxies.
@@ 4.1 CASTOR survey images
Here we present example CASTOR images created using our image simulator, for each of the planned CASTOR surveys: Wide, Deep, and Ultra-Deep. For each exposure time and filter combination, we create six images of various CASTOR FOVs within the 2-deg2 SAM catalogue. This allows us to test our codes and predictions with more galaxies, as well as giving estimates for the effect of cosmic variance on the number of galaxies expected in each image. In this section we show only one of these FOVs as an example, although we use all of these in our quantitative analysis in Section 5. We note that our images are made only using a galaxy catalogue, and include no foreground stars. Thus we cannot investigate issues such as star–galaxy blending, where some of the detectable galaxies in our images would, in reality, be obscured and undetectable because of overlapping foreground stars. However, this allows us to more easily extract sources and perform a comparison to our input galaxy catalogue, which is adequate for our galaxy-focused study. With the high resolution of CASTOR, we expect that this will have a negligible impact on the completeness limits that we estimate in this work.
@@ 4.2 Comparison to GALEX and Rubin
To simulate GALEX images for a comparison to CASTOR, we use its telescope parameters diameter m, pixel scale /pixel, a gain of 1.0 e-/ADU, with no dark current or read noise. We simulate images in the NUV filter, assuming a PSF FWHM of , which we assume is a Gaussian in shape (Morrissey et al., 2005), and a sky background level of 26.5 mag/arcsec2 (Martin et al., 2005). For exposure times, we use those from the All-sky Imaging Survey (AIS), 0.1ks, Medium Imaging Survey (MIS), 1.5ks, Deep Imaging Survey (DIS), 30ks, and the Ultra-Deep Imaging Survey (UDIS), 200ks (Martin et al., 2005). The resulting images are shown in Figure 3. These clearly show the improvement in UV-imaging that CASTOR will have over GALEX, with significantly improved depth and resolution. The GALEX resolution was in the NUV, compared to CASTOR’s planned . This 30-fold improvement in resolution will mitigate the problem of source blending in GALEX imaging, and significantly improve our ability to identify UV sources and accurately measure their properties — both in the local universe and across cosmic timescales (Martin et al., 2005; Noeske et al., 2007; Salim et al., 2007; Gil de Paz et al., 2007; Thilker et al., 2007).
@@
To make simulated Rubin images, we use a diameter m, a pixel scale of /pixel, dark current of 0.02 e-/pixel/s, a read noise of 9 e-/pixel, and a gain of 1.0 e-/ADU (O’Connor, 2019). We simulate images in the g-band, with a sky background of 22.26 mag/arcsec2 (Ivezić et al., 2019). We assume that the PSF is the convolution of two Gausssians, one with FWHM and one with , each with the same flux, to be consistent with Imsim, the LSST simulation software package111https://lsstdesc.org/imSim. We simulate exposure times of 30s, corresponding to a single visit, 300s, the 1 year depth, 3000s, the 10 year depth, and 39000s, the final depth in each of LSST’s four Deep Drilling Fields (see Côté et al. 2024, in preparation). The resulting images are shown in Figure 3. While the improvement of the CASTOR images is not as dramatic as for GALEX, the zoomed-in panels still clearly show the effect of the increased resolution of CASTOR over LSST. With LSST’s FWHM due to atmospheric seeing, the space-based CASTOR offers a factor of improvement in spatial resolution.